40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Feb 3 04:01:55 2015, diego, Update, LSC, CM servo & AO path status CARM_03-02-2015_031754.pdf
    Reply  Tue Feb 3 16:36:33 2015, ericq, Update, LSC, CM servo & AO path status yarmAO.pdf
       Reply  Mon Feb 9 17:47:17 2015, diego, Update, LSC, CM servo & AO path status 
Message ID: 10969     Entry time: Tue Feb 3 16:36:33 2015     In reply to: 10966     Reply to this: 10991
Author: ericq 
Type: Update 
Category: LSC 
Subject: CM servo & AO path status 

I have removed REFLDC and the SR560 offsetter from the CM board IN2. Now, analog AS55 I lives there, for our single arm testing. (Analog I has more of the single arm Y PDH signal in it). REFL11 has been reconnected to IN1. 


With ITMX super misaligned, Diego and I locked the Y-arm with the AO path on AS55, ultimately at 4kHz bandwidth, but with plenty of gain margin. We didn't allocate the gains too intelligently, and had the CM board input gain slider maxed out, but plenty of headroom in the digital and AO sliders, making it inconvenient to up the UGF even more, to engage the super boosts. However, since this is just a test case to make sure we still can AO lock, I'm not too worried about this. 

Since LSC FMs and such had changed around, old recipies didn't neccesarily work 1:1. Diego is writing a script for the current recipe, and will post an elog with the steps. 

Gains and signs are able to be tracked by loop TFs, the real sticking point is a stable crossover. We used the 1.6k:80 hardware filter in the CM board to give the AO Path a 1/f shape in the crossover region, and undid it digitally in the CM_SLOW input FM. However, we do use a 300:80 in the MC2 sus FM to make the digital loop like 1/f^2 around the crossover, once a little bit of AO has come in to pull up the digital loop's phase. We used the CARM filter bank to do this, so I think we should be able to use a similar technique to do it in the PRFPMI case, as long as the coupled cavity pole is around ~100Hz. 

Attached are a few OLTFs from the progression.

Attachment 1: yarmAO.pdf  43 kB  | Hide | Hide all
yarmAO.pdf
ELOG V3.1.3-